OpenStudy (anonymous):

Evaluate the integral of tan 3x dx

OpenStudy (blurbendy):

use substitution. let u = 3x du=3dx dx = 1/3du 1/3 (integral sign) tan(u)du = - 1/3ln|cos(u)| + C = -1/3ln|cos(3x)|

OpenStudy (blurbendy):

+ C